# Groups and quandles

### v1

@@ Abstract.
We are intereseted in quandles and their enveloping groups. Various results are proven. We show that a quandle $Q$ and its image in the enveloping group $G(Q)$ have isomorphic enveloping groups. The image quandle is injective. For $Q$ a finite quandle, we show that $G(Q)$ admits a faithfull representation $\rho:G(Q)\to GL_{n}({\mathbb{Z}})$ for some $n$; an irreducible representation of $G(Q)$ over $\mathbb{C}$} is finite dimensional an its degree divides the order of the group $Inn(Q)$ of inner automorphism of $Q$. We determine the Malcev Lie algebra and the rational cohomology ring of $G(Q)$ for $Q$ finite. We prove that a finite injective quandle is a subquandle (for conjugacy) of a finite group. We also prove that the only finite subquandles (for conjugacy) of uniquely divisible groups are trivial quandles and that morphisms from quandles to nilpotent groups (for conjugacy) are constant on the indecomposable components. Implication of these results are considered.
@@
We prove in section 4, that for $Q$ finite $G(Q)$ admits a faithfull representation $\rho:G(Q)\to GL_{n}({\mathbb{Z}})$, for some $n$. This proves that every finite injective quandle admits an injective representation $\rho:Q\to Conj(GL_{n}({\mathbb{Z}}))$ for some $n$. The linearity of $G(Q)$ for $Q$ finite implies that $Q$ is a subquandle for conjugacy of some finite group. We also show that irreducible representations of a finite quandle $Q$ over $\mathbb{C}$} (equivalently irreducible representations of $G(Q)$) are finite dimensional and that the degree of such representations divides the order of the group $Inn(Q)$ of inner automorphisms of $Q$.
@@ 3. Injective quandles and an intermidiate functor
One can wonder wether $\varphi_{Q}:Q\to G(Q)$ is always injective. The answer is no. We will give an example and a general argument.
@@ 4. Enveloping group of finite quandles and representations
For $Q$ a quandle it follows from the definition of a quandle we took that for $x\in X$ the map $L_{x}:Q\to Q$ given by $L_{x}(y)=x\triangleright y$ is a bijective quandle morphism. The group $Inn(Q)$ is the subgroup of the group of bijections of $Q$ generated by the maps $L_{x}$ for $x\in X$.
@@
$V$ is irreducible. Hence, any vector is cyclic and since $G$ is countable $V$ has countable dimension over $\mathbb{C}$}. The commutant $D$ of the representation has hence a countable dimension over $\mathbb{C}$}. Indeed, the linear morphism $D\to V,T\mapsto T(v)$ for $v\in V$, $v\neq 0$ is injective by Schur’s lemma. By Schurs lemma $D$ is a also a division algebra over $\mathbb{C}$} (the scalar matrices). Since the only division algebra of countable dimension over $\mathbb{C}$} is $\mathbb{C}$} itself, $D$ is reduced to the scalar matrices and the lemma follows. ∎
Theorem 4.8.
Let $G$ be a countable group with center of finite index. The irreducible representations of $G$ over $\mathbb{C}$} are finite dimensional.
@@
Let $\theta:G\to GL(V)$ be an irreducible representation. Denote by $\mathcal{Z}$ the center of $G$. Let $s:G/\mathcal{Z}\to G$ be a section and define $\gamma:G/\mathcal{Z}\times G/\mathcal{Z}\to\mathcal{Z},(g,h)\mapsto s(g)s(h)s(gh)^{-1}$. By the previous lemma the center acts by scalar hence we have a 2-cocyle $\alpha:G/\mathcal{Z}\times G/\mathcal{Z}\to{\mathbb{C}}^{*}$ mapping $(g,h)$ to the scalar correponding to the action of $\gamma(g,h)$ on $V$. Now let $\rho:{\mathbb{C}}[G/\mathcal{Z}]\to GL(V)$ be the linear map defined by $\rho(g)=\theta(s(g))$ for $g\in G/\mathcal{Z}$. One has that :
$\alpha(g,h)\rho(gh)=\rho(g)\rho(h),$
for $g,h\in G/\mathcal{Z}$. This proves that $\rho$ is a represenation of the twisted algebra (${\mathbb{C}}[G/\mathcal{Z}],\star)$ where $\star$ is the associataive product (since $\alpha$ is a $2$-cocyle) given by :
$g\star h=\alpha(g,h)gh,$
for $g,h\in G/\mathcal{Z}$. The representation $\rho$ is irreducible since $\theta$ is irreducible. Since $\mathcal{Z}$ has finite index ${\mathbb{C}}[G/\mathcal{Z}]$ is finite dimensional and hence $V$ is finite dimensional. Indeed, irreducible representations of a finite dimensional algebra are quotients of the algebra by a maximal left ideal (one can use a cylic vector for instance to construct a morphism form the algebra to the vector space of the representation). ∎
@@
For $Q$ a finite quandle, the irreducible representations of $Q$ or equivalently of $G(Q)$ are finite dimensional.
@@
follows from the last theorem, the fact that $G(Q)$ for finite quandles is finitely generated and corollary 4.2. ∎

### v2

@@ Abstract.
We are intereseted in quandles and their enveloping groups. Various results are proven. We show that a quandle $Q$ and its image in the enveloping group $G(Q)$ have isomorphic enveloping groups. The image quandle is injective. For $Q$ a finite quandle, we show that $G(Q)$ admits a faithfull representation $\rho:G(Q)\to GL_{n}({\mathbb{Z}})$ for some $n$; an irreducible representation of $G(Q)$ over $\mathbb{C}$} is finite dimensional an its degree divides the order of the group $Inn(Q)$ of inner automorphism of $Q$ and is bounded by $\sqrt{|Inn(Q)|}$. We determine the Malcev Lie algebra and the rational cohomology ring of $G(Q)$ for $Q$ finite. We prove that a finite injective quandle is a subquandle (for conjugacy) of a finite group. We also prove that the only finite subquandles (for conjugacy) of uniquely divisible groups are trivial quandles and that morphisms from quandles to nilpotent groups (for conjugacy) are constant on the indecomposable components. Implication of these results are considered.
@@
We prove in section 4, that for $Q$ finite $G(Q)$ admits a faithfull representation $\rho:G(Q)\to GL_{n}({\mathbb{Z}})$, for some $n$. This proves that every finite injective quandle admits an injective representation $\rho:Q\to Conj(GL_{n}({\mathbb{Z}}))$ for some $n$. The linearity of $G(Q)$ for $Q$ finite implies that $Q$ is a subquandle for conjugacy of some finite group. We also show that irreducible representations of a finite quandle $Q$ over $\mathbb{C}$} (equivalently irreducible representations of $G(Q)$) are finite dimensional and that the degree of such representations divides the order of the group $Inn(Q)$ of inner automorphisms of $Q$; moreover, the degree is bounded by $\sqrt{|Inn(Q)|}$.
@@ 3. Injective quandles and an intermidiate functor
One can wonder wether $\varphi_{Q}:Q\to G(Q)$ is always injective. The answer is known ([JO82]) to be no. We will give an example from [JO82] and a general argument.
@@ 4. Enveloping group of finite quandles and representations
For $Q$ a quandle it follows from the definition of a quandle we took that for $x\in Q$ the map $L_{x}:Q\to Q$ given by $L_{x}(y)=x\triangleright y$ is a bijective quandle morphism. The group $Inn(Q)$ is the subgroup of the group of bijections of $Q$ generated by the maps $L_{x}$ for $x\in Q$.
@@
$V$ is irreducible. Hence, any vector is cyclic and since $G$ is countable $V$ has countable dimension over $\mathbb{C}$}. Therefore, the lemma follows from Dixmier’s lemma.∎
Proposition 4.8.
Let $G$ be a countable group with center $\mathcal{Z}(G)$ of finite index. An irreducible representation of $G$ over $\mathbb{C}$} is finite dimensional and the degree of such represenation is bounded by $\sqrt{|G/\mathcal{Z}(G)|}$.
@@
Let $\theta:G\to GL(V)$ be an irreducible representation of a group as in the proposition. Let $s:G/\mathcal{Z}(G)\to G$ be a section and $v$ be a cyclic vector of $V$. By the previous lemma the center acts by multiplication by scalars and hence the span of the vectors $gv$ for $g\in G$ is equal to the span of the vectors $s(h)v$ for $h\in G/\mathcal{Z}(G)$. Since the center has finite index the representation is finite dimensional. Let $\rho:{\mathbb{C}}[G]\to End(V)$ be the algebra morphism corresponding to the irreducible representation $\theta$ of $G$. Since $V$ is finite dimensional by the density theorem $\rho$ is surjective. On the other hand, since the center acts by scalars, the image of $\rho$ is linearly generated by $\rho(s(h))$ for $h\in G/\mathcal{Z}(G)$. Hence, $dim(V)^{2}\leq|G/\mathcal{Z}(G)|$. We have proved the proposition. ∎
@@
For $Q$ a finite quandle, an irreducible representations of $Q$ or equivalently of $G(Q)$ is finite dimensional; the degree of such representation is bounded by $\sqrt{|Inn(Q)|}$.
@@
follows from the last theorem, the fact that $G(Q)$ for finite quandles is finitely generated, corollary 4.2 and the fact (that we will see later) that we have a surjective morphism from $Inn(Q)\to Inn(Q_{conj})$. ∎
@@
Remark 4.12.
A finite quandle have infinitely many irreducible representation. For instance, as we will see the abelianisation of $G(Q)$ is a free abelian group and hence $G(Q)$ and $Q$ have infinitely many non equivalent $1$-dimensional representation.
Remark 4.13.
The group $G(Q)$ is of type I, since it is countable and contains an abelian subgroup of finite index, its center.